OpenStudy (anonymous):

How much does a watch lose per day, if its hands coincide every 64 minutes?

OpenStudy (shubhamsrg):

you mean to ask how much times they collide in 24hrs ?

OpenStudy (shubhamsrg):

how many times **

OpenStudy (anonymous):

i mean loss in 24 hrs

OpenStudy (shubhamsrg):

loss of what ?

OpenStudy (anonymous):

time

OpenStudy (shubhamsrg):

and how much time is lost per collision ?

OpenStudy (anonymous):

yes

OpenStudy (shubhamsrg):

what yes? that should be given in the question..bah..let it be anyways,,from the given info,,you can only find how many times they collide in 24hrs no. of mins in 24 hrs. = 24 *60 thus,,no of collisions = 24*60/64 = 180/8 = 22.5 i.e. 22 complete collisions